OpenStudy (anonymous):

Solve 2x2 + 8x - 12 = 0 by completing the square. What are the solutions

OpenStudy (amistre64):

we gotta divide out the 2 then and we are left with: x^2 +4x -6 to complete a square means this: \begin{array}c x^2&x&x\\ x&1&2\\ x&3&4 \end{array} x^2 + 4x + 4 is a complete square; so add 4 and subtract 4 to our equaiton

OpenStudy (dumbcow):

factor out a 2 x^2 +4x -6 = 0 take half middle coefficient and square it x^2 +4x +4 -6 = 4 (x+2)^2 = 10 x+2 = sqrt(10) x = +-sqrt(10) -2

OpenStudy (amistre64):

x^2 +4x +(4 -4) -6 = 0 ; regroup (x^2 +4x +4) +(-4 -6) = 0 ; compact the square (x+2)^2 -10 = 0 (x+2)^2 = 10 ; now this would be ideal if we had (sqrt(10))^2 but we have to subtract a 2 out if to make that work. x = -2 +- sqrt(10)